# inkscapelatex=false

### v2

@@ 3.3 Place Field Analysis
In order to understand how the agent executes a successful strategy and whether the representations in the hippocampus-inspired model would also align with the known physiology of place cells, we further computed activity maps of the units for the different layers of the model network. For the input units (DG), we (by construction) observe sparse activation that is distributed across few locations in the map (Fig. 5). More interestingly, learning the weights from the visual encoder to DG leads to orthogonalization, such that DG population activity over time develops a unique representation of individual locations (Fig. 6). Units positioned later in CA3 sequences (e.g. columns no.48 and no. 64 in Fig. 5) show broader spatial tuning, in line with experimental reports on the CA3 spatial code being less sparse than in DG (Jung and McNaughton, 1993). In contrast, the input to LSTM agents did not show orthogonalization of the learned input projection (Fig. A11) and the spatial tuning of the LSTM units are disperse across the entire map, distinct from localized place-cell-like tunings (Fig. A9.
Figure 7: SI of units in CA3 as a function of their order in a sequence and its evolution across epochs. Each sequence is indicated by a unique color across epochs.
@@ 3.4 Spatial Information Analysis
Qualitative inspection of the spatial tunings in Fig. 5 is also quantitatively supported by the distributions of spatial information (SI) of the neural activity maps: CA3 units activated $\sim 16$ steps after the input have larger SI (Fig. 7). This corroborates our training results that agents with sequence length $L=16$ are able to acquire the task reasonably well (Fig. 3). Units activated later in a sequence exhibit smaller spatial information (Fig. 7).
@@ 3.5 Population-level Representations
We next examined population-level representations using mean population vector (PV) correlations, defined as the Pearson correlation between two activity vectors separated by a spatial displacement $\Delta x,\Delta y$. This measure can be interpreted as a spatial kernel learned by the network (Fig. 8). Kernels from all layers (DG, CA3, and Decoder layers) exhibit some dependence on distance throughout learning. However, CA3 kernels are smoother than the DG kernels. Both DG and CA3 show lower correlation values compared to the actor–critic layers. Notably, the first Decoder layer (Decoder1, biologically approximately equivalent to area CA1) develops the most pronounced spatial tuning (Fig. 8), consistent with the single-unit selectivity observed in Fig. A6. After the reward location was changed (“new”), kernels became less sharply defined, indicating that the representations are disrupted although behavior is adapted. Most layers in the agent with LSTM core and dense input did not show spatial kernels with strong displacement dependency. Only the LSTM output units showed a gradually refined spatial kernel during learning but it is strongly non-isotropic (Fig. A12).
Figure 8: Kernel representation of distance. A Mean Pearson correlation (color code) between two population activity vectors as a function of their spatial displacement. Rows correspond to different layers (brain regions) of the network, columns to learning stages (epochs). B Left: summary of population vector correlation as a function of distance. Right: summary of population vector correlation as a function of orientation. Radial coordinate range: (-0.5,1.1).
Over training, a stable place code gradually emerges (Fig. A7), with DG similarities amplified in downstream layers. Interestingly, the similarity between fully trained networks (epoch 6) and the novel reward condition (“new”) is higher than between the naive (epoch 0) and trained (epoch 6) networks, suggesting that the agent acquires generalizable knowledge about the arena’s spatial layout.
@@ 4 Discussion
In this study, we presented a minimal model of the hippocampus that enables navigation in a vision-based virtual environment and reproduces key phenomena of spatial representations observed in the mammalian hippocampus.
The cornerstone of our model—the sequential connectivity in CA3—is biologically plausible and inspired by established findings on theta sequence firing Dragoi and Buzsáki (2006); Foster and Wilson (2007). By constraining CA3 recurrent connectivity as a dynamical reservoir and limiting training to input and output weights via reinforcement learning, we clearly isolate the effect of this theta sequence generator.
@@
Analysis of neural activities revealed clear parallels with experimental observations of place cell properties, including robust place field formation (Fig. 5), population-level encoding of spatial distance (Fig.8), progressive orthogonalization of DG outputs (Fig. 6), and dynamic remapping triggered by changes in reward locations (Fig.A7), consistent with experimental findings Leutgeb and Leutgeb (2007); Fenton (2024).
@@
The CA3 module expands sparse DG codes into a temporally smoothed canonical basis set, providing long-horizon history without the indiscriminate feature mixing of fully connected RNNs. This is especially beneficial under sparse input, where immediate sensory signals are limited and long-range context is critical for policy learning.
At first sight, sparsifying input to a recurrent core seems purely detrimental—leaving the agent “blind” most of the time. Yet it also filters out noisy, non-informative cues, making each supra-threshold input more reliably tuned to restricted regions of space. Through sequence propagation and policy learning, representations of other locations can then be inferred and stabilized. This mechanism is consistent with habitual trajectories: neuronal activity late in a sequence only remains informative if the policy converges onto consistent paths.
In contrast, with dense input streams, conventional recurrent cores such as LSTMs performed better, consistent with their ability to rapidly integrate continuous signals. Our CA3 model, however, performed worse under dense input. A likely reason is that its long-range memory buffer provides little benefit when input is already abundant, and temporal smoothing obscures informative short-term variations.
@@
We proposed a parsimonious account of hippocampal theta sequences: they can be intrinsically maintained without requiring external input. While actual theta sequences in biological systems likely arise from multiple factors and vary across contexts Chance (2012); Schlesiger et al. (2015); Ahmadi et al. (2025), our abstraction nonetheless reproduced hallmark hippocampal phenomena, including robust place fields (Fig. 5), progressive DG orthogonalization (Fig. 6), distance-dependent population kernels (Fig. 8), remapping after goal changes, and, by construction, theta sequences (Foster and Wilson, 2007). These representational effects were linked to performance and align with evidence on sparse DG activity and CA3 sequences (Leutgeb and Leutgeb, 2007; Fenton, 2024).
@@
The comparison between our hippocampus-inspired model and LSTM agent with dense input demonstrate the distinction of learned strategies and representations, even though the performance indicators are comparable (Figs. A12A11A8A9A10).
@@
Our model predicts that larger environments or sparser inputs require longer sequences for successful navigation, consistent with adaptability of sequence length (Wikenheiser and Redish, 2015; Farooq and Dragoi, 2019). More broadly, it suggests that hippocampal spatial representations could rely largely on intrinsic sequence-generating circuitry, with experience primarily shaping feedforward and readout connections. It also offers an explanation for how place cells can persist despite lesions of entorhinal cortex (Brun et al., 2008; Steffenach et al., 2002; Schlesiger et al., 2015).
@@
Several extensions follow naturally. On the biological side, incorporating local plasticity rules into the DG–CA3 pathway or CA3 readouts would align the model more closely with known mechanisms while preserving the benefits of prewired dynamics. Interactions with entorhinal inputs and path integration could further clarify their complementary roles in spatial cognition. Moreover, theta sequences have also been observed in other brain areas, raising the possibility of hierarchical coordination across regions.
@@
On the ML side, combining sparse sequential reservoirs with learned SSMs may yield hybrid architectures capable of adapting flexibly across input regimes. More broadly, a bottom-up approach like ours—focusing on circuit motifs rather than top-down expert informed objectives—suggests a path toward scaling to larger networks and more complex tasks.
@@
Beyond ML implications, our framework illustrates how modern RL can serve as a testbed for neuroscience theories. By embedding a biologically inspired model within a reinforcement-learning agent, we can study brain computation beyond intermediate objectives such as sensory processing, cognition, or motor control, and consider the entire perception–action cycle—particularly useful for association cortices such as hippocampus and prefrontal cortex where representations reflect both the input and action planning.
Furthermore, the behavioral and representational comparison between CA3 agent and LSTM agent showed a useful direction of research - qualitatively distinct strategies beyond performance benchmarking can provide useful insights for architectural designs.
@@ Conclusion
A minimal, sparsely driven sequence generator in an actor-critic agent not only is sufficient for successful navigation but also for hippocampus-like spatial representations.
@@ 5 Limitations
Our results depend on the environment specifications and hyperparameter settings. We chose a reasonably difficult environment setting where the difference between architectures can be highlighted. It does not imply the effect can generalize to all environments. We tested the agent with L=64 and R=8 on 5 other randomly generated maps, it could converge at a similar speed with the map shown in other results A2. We also tested different frame-skip number, larger frame-skip leads to faster learning, and the effect of sequence length on performance is weaker (Fig. A13). This is probably due to the fact that our environment has sparse obstacles and that larger frame-skip leads to effectively longer sequence and memory. Most results in our report use 4 policies for population based training and 8 environments per worker. These parameters also change the dynamics of learning. having more policies made agents with L=64 learn slower and L=8 learn slower. Further systematic investigation of these factors is required to establish broader applicability.
@@
6 Implementation Details
6.1 Environment
@@
Since we are mainly interested in modeling the way-finding aspect of navigation [Tolman, 1948, Kuipers, 2000], the MLP also receives average depth of the pixels, down-sampled to 10 horizontal pixels to aid motor control (going around walls and avoiding collisions). This procedure is supposed to mimic the sensory motor collision avoidance habits that do not depend on hippocampus. This depth information is directly routed to the decoder layer, so the representations in the recurrent core and its input are not directly influenced.
6.2 Architecture
Table A3: Network architecture of the CA3 agent. The ResNet encoder is pretrained and frozen during reinforcement learning to isolate the contribution of hippocampal modules.
Component Specification
Visual encoder ResNet using the default parameters from sample factory [Petrenko et al., 2020]
3 convolutional stages (3→16→32→32 channels) each followed by max pooling;
6 residual blocks total (two per stage), each with [ReLU → Conv(3×3) → ReLU → Conv(3×3)] + skip connection;
final ReLU applied to output feature map.
Flatten + MLP head Flatten → Linear(3456→256) → ReLU; output = 256-D embedding.
DG projection Linear(256→16) → BatchNorm1d(momentum=0.05, affine=False) → ReLU; intercept threshold = 2.43.
CA3 (recurrent core) Fixed linear shift-register reservoir; input features = 16, sequence length $L=64$, repetition $R=8$;
total state size $16\times(64+8-1)=1136$.
Decoder MLP: Linear(1136+10→128) → ReLU → Linear(128→128) → ReLU.
Critic head Linear(128→1).
Actor head Linear(128→5) (softmax over 5 discrete actions).
Training scheme Asynchronous PPO [Espeholt et al., 2018, Petrenko et al., 2020] ;
ResNet weights frozen; DG, Decoder, Actor, and Critic trained.
CA3 agent has learnable parameters in DG projection and from CA3 to Decoder, the CA3 output in the full version has 1136*128=149504 parameters. LSTM has learnable parameters $4*(m^{2}+mn)$ where m is hidden size and n is input and output size. Solving this gives a hidden size of roughly 137, which we used in the implementation. SSM agents have matching state size with CA3 agent. They also have block-diagonal recurrent weights, where the weights in each block are obtained through the original implementation by Gu et al. [2020] with zero-order-hold discretizations.
6.3 Actor Critic
We optimize an actor–critic objective with advantage estimates based on the clipped PPO formulation [Schulman et al., 2017]. The total loss combines a clipped policy surrogate, a value regression term, and an entropy regularizer:

### v3

@@ 3.3 Place Field Analysis
In order to understand how the agent executes a successful strategy and whether the representations in the hippocampus-inspired model would also align with the known physiology of place cells, we further computed activity maps of the units for the different layers of the model network. For the input units (DG), we (by construction) observe sparse activation that is distributed across few locations in the map that is further refined in the subsequent linear readout in decoder layer 1 (Fig. 4AB). Hidden units in the LSTM agent show non-localized tuning, distinct from typical place cells (Fig. 4C). More interestingly, learning the weights from the visual encoder to DG leads to orthogonalization, such that DG population activity over time develops a unique representation of individual locations (Fig. 5A). Units farther away from DG input in CA3 sequences show broader spatial tuning (Fig. 4A; quantified in Fig. 5B) matching previous reports (Jung & McNaughton, 1993; Parra-Barrero et al., 2021). The place fields of DG and CA3 units gradually stabilize through learning and exhibit remapping after adapting to a new reward location, measured by shifts of the place field center of mass (Fig. A10). In contrast, the input to LSTM agents did not show orthogonalization of the learned input projection (Fig. A14).
@@ 3.4 Spatial Information Analysis
Qualitative inspection of the spatial tunings in Fig. 4 is also quantitatively supported by the distributions of spatial information (SI) of the neural activity maps: CA3 units activated $\sim 16$ steps after the input have larger SI (Fig. 5C). This corroborates our training results that agents with sequence length $L=16$ are able to acquire the task reasonably well (Fig. 2). Units activated later in a sequence exhibit smaller spatial information (Fig. 5C).
@@ 3.5 Population-level Representations
We next examined population-level representations. We employ a novel method inspired by population vector (PV) correlation and representational (dis)similarity analysis (Kornblith et al., 2019; Kriegeskorte & Wei, 2021). It measures the mean population vector correlation in a pair of location bins grouped by the spatial displacement $\Delta x,\Delta y$. This measure can be interpreted as a spatial kernel learned by the network (Fig. 6). An unbiased representation of spatial geometry would show a kernel function that is isotropic, and smoothly and monotonically depends on the distance between locations bins. Conversely, the representation could be restricted to specific location bin pairs, displacement along specific orientation, specific spatial frequency, or simple visual resemblance. Kernels from all layers (DG, CA3, and Decoder layers) exhibit some dependence on distance throughout learning. However, CA3 kernels are smoother than the DG kernels. Both DG and CA3 show lower correlation values compared to the decoder layer. Notably, the Decoder layer 1 develops the most pronounced spatial tuning (Fig. 6A), consistent with the single-unit selectivity observed in Fig. A8. After the reward location was changed (“new”), kernels became less sharply defined, indicating that the representations are disrupted although behavior is adapted. Most layers in the agent with LSTM core and dense input did not show spatial kernels with strong displacement dependency. Only the LSTM output units showed a gradually refined spatial kernel during learning but it is strongly non-isotropic (Fig. 6B; Fig. A15). The dependence of representational similarity on distance and orientation is quantified in Fig. 6C.
Figure 6: Kernel representation of distance. A Mean Pearson correlation (color code) of pairs of population activity vectors (PV) as a function of spatial displacement. Rows correspond to different layers (brain regions) of the network, columns to learning stages (epochs). B Same as A, but in dense LSTM agents. C PV correlation as a function of distance (left) and orientation (right).
Over training, a stable place code gradually emerges and remaps after learning the new reward location, indicated by the correlation of population vectors across training epochs (Fig. A9). Interestingly, the similarity between fully trained networks (epoch 6) and the novel reward condition (“new”) is higher than between the naive (epoch 0) and trained (epoch 6) networks, suggesting that the agent acquires generalizable knowledge about the arena’s spatial layout.
@@ 4 Discussion
In this study, we presented a minimal model of the hippocampus that enables navigation in a vision-based virtual environment and reproduces key phenomena of spatial representation observed in the mammalian hippocampus.
The cornerstone of our model — the sequential connectivity in CA3 — is biologically plausible and inspired by established findings on theta sequence firing (Dragoi & Buzsáki, 2006; Foster & Wilson, 2007; Yiu et al., 2022). By constraining CA3 recurrent connectivity as a dynamical reservoir and limiting training to input and output weights via reinforcement learning, we clearly isolate the effect of this theta sequence generator.
@@
Analysis of neural activities revealed clear parallels with experimental observations of place cell properties, including robust place field formation (Fig. 4), population-level encoding of spatial distance (Fig. 6), progressive orthogonalization of DG outputs (Fig. 5A), and dynamic remapping triggered by changes in reward locations (Fig. A9 A10), consistent with experimental findings (Leutgeb & Leutgeb, 2007; Fenton, 2024).
@@
The CA3 module expands sparse DG codes into a temporally smoothed canonical basis set, providing long-horizon history without the indiscriminate feature mixing of fully connected RNNs. This is especially beneficial under sparse input, where immediate sensory signals are limited and long-range context is critical for policy learning. The repetition parameter $R$ functions as a built-in prior of the temporal smoothness of latent states, thereby smoothing the CA3 spatial tuning via the agent’s movement, at the cost of imprecise timing of input memory.
At first sight, sparsifying input to a recurrent core seems purely detrimental—leaving the agent “blind” most of the time. Yet it also filters out noisy, non-informative cues, making each supra-threshold input more reliably tuned to restricted regions of space. Through sequence propagation and policy learning, a spatially smooth representations can then be stabilized. This mechanism is consistent with habitual trajectories: neuronal activity late in a sequence only remains informative if the policy converges onto consistent paths.
Under dense input, LSTMs perform better, likely because the abundant weak signals can be accumulated to inform space, reducing the need for long-horizon buffering of strong signals. These intuitions are supported by comparative experiment where different levels of Gaussian noise were added to the pixel input to suppress weak signals. The DG+CA3 agent suffers less from increasing noise level than the LSTM agent (Fig. A5; Tab. A4).
@@
We proposed a parsimonious account of hippocampal theta sequences: they can be intrinsically maintained without requiring external input. While actual theta sequences in biological systems likely arise from multiple factors and vary across contexts (Chance, 2012; Schlesiger et al., 2015; Yiu & Leibold, 2023; Ahmadi et al., 2025), our abstraction nonetheless reproduced hallmark hippocampal phenomena, including robust place fields (Fig. 4), progressive DG orthogonalization (Fig. 5A), distance-dependent population kernels (Fig. 6), remapping after goal changes (Fig. A9 and Fig. A10), and, by construction, theta sequences (Foster & Wilson, 2007). These representational effects were linked to performance and align with evidence on sparse DG activity and CA3 sequences (Leutgeb & Leutgeb, 2007; Fenton, 2024).
Importantly, the visual encoder can be viewed as capturing the visually driven components of entorhinal cortex (EC) input to DG and CA3: the dorsal “where” and ventral “what” streams that project to medial and lateral EC (Wang et al., 2011). The depth sensor signal fed to Decoder can be seen as temporoamonic pathway from EC to CA1. To reflect the multi-modal nature of EC, additional modalities could be incorporated as input features.
@@
The comparison between our hippocampus-inspired model and an LSTM agent with dense input demonstrate the distinction of learned strategies and representations, even though the performance indicators are comparable (Figs. A15A14A11A12A13).
@@
Our model predicts that larger environments or sparser inputs require longer sequences for successful navigation, consistent with developmental adaptability of sequence length (Wikenheiser & Redish, 2015; Farooq & Dragoi, 2019). More broadly, it suggests that hippocampal spatial representations could rely largely on intrinsic sequence-generating circuitry, with experience primarily shaping feedforward and readout connections. It offers an explanation for how place cells can persist despite lesions of entorhinal cortex (Hales et al., 2014; Schlesiger et al., 2015). The parsimonious nature of our model also provides a unified mechanism how navigation can be built upon sequences, no matter how they arise and how they differ across species. This is particularly important in species without prominent theta oscillations, e.g. bats show hippocampal sequences locked to wingbeats (Forli et al., 2025).
@@
Several extensions follow naturally. On the biological side, incorporating local plasticity rules into the DG–CA3 pathway or CA3 readouts would align the model more closely with known mechanisms while preserving the benefits of prewired dynamics. Interactions with path integration in medial EC could further clarify their complementary roles in spatial cognition. Moreover, theta sequences have also been observed in other brain areas, raising the possibility of hierarchical coordination across regions.
@@
On the ML side, combining sparse sequential reservoirs with learned SSMs may yield hybrid architectures capable of adapting flexibly across input regimes. More broadly, the bottom-up approach—focusing on circuit motifs rather than top-down expert informed objectives—suggests a path toward scaling to larger networks and more complex tasks.
@@
Our framework showcases how modern RL can serve as a testbed for computational neuroscience. By embedding a biologically inspired hippocampal circuit within an end-to-end reinforcement-learning agent, the model enables us to study brain computation in the context of the full perception–action loop, rather than in isolated modules for sensory processing, cognition, or motor control. Our approach thereby relates to ongoing discussions in machine learning regarding the “Bitter Lesson”, which cautions against relying on handcrafted intermediate representations in favor of scalable end-to-end optimization. Importantly, our structural priors (sparseness, sequences) do not prescribe the representational solution; place fields, distance-dependent kernels, and remapping emerge through end-to-end learning. In this sense, physiological constraints act as an inductive bias that narrows the hypothesis space without sacrificing scalability, providing a middle ground between unconstrained black-box models and heavily engineered representational schemes. This perspective – imposing minimal structural constraints, such as sparse connectivity, rather than making apriori assumptions – is particularly relevant for association cortices such as hippocampus and prefrontal cortex, where emerging representational patterns are shaped jointly by sensory inputs, movement statistics, and behavioral goals (Lin et al., 2023).
An important conceptual point concerns how our representations relate to existing theoretical frameworks. Our CA3 representations relate to both successor-representation (SR) and reward/value-based frameworks, but with important distinctions. As in SRs, CA3 activity is policy dependent and shows a consistent temporal ordering: because the CA3 module propagates an intrinsic sequence, each unit predicts its downstream neighbors. This produces SR-like anticipatory structure, yet it arises solely from the fixed physiology-inspired architecture—CA3 does not learn a predictive map and receives no TD updates or discounted occupancy signal.
At the same time, unlike many previous RL studies of hippocampus (Kumar et al., 2022; Leibold, 2020), that takes allocentric input and analyze how policy shapes and utilizes them, our contribution is complementary: we show how localized, Gaussian-like place tuning can emerge from egocentric observations when combined with sparse DG input and intrinsic CA3 sequences. Our work also distinguishes itself from many other work that takes egocentric input but where the spatial similarity structure is already reflected in the input (e.g. small room where each wall has a different color; Vijayabaskaran & Cheng (2022); Raju et al. (2024)). Our environment is large maze, with uniform looking obstacles scattered in the arena, accounting for a more realistic discrepancy between visual representation and spatial representation.
The behavioral and representational divergence between the CA3-agent and the LSTM-agent illustrates a promising direction for future comparative work. Differences in qualitative strategies—not only performance—can reveal how architectural constraints shape navigation, representational geometry, and generalization, providing insights that are difficult to obtain from performance benchmarks alone.
@@ Conclusion
A minimal, sparsely driven sequence generator in actor-critic agent not only supports successful navigation but also gives rise to hippocampus-like spatial representations.
5 Acknowledgements
This work was funded by the German federal ministry for research, technology and space travel (BMFTR) under grant number 01GQ2510 and the BrainLinks-BrainTools Center.
@@ Appendix A Limitations
Our results depend on the environment specifications and hyperparameter settings. We chose a reasonably difficult environment setting where the difference between architectures can be highlighted. It does not imply the effect can generalize to all environments. We tested the agent with L=64 and R=8 on 5 other randomly generated maps, it could converge at a similar speed with the map shown in other results (Fig. A3). We also tested different frame-skip number, larger frame-skip leads to faster learning, and the effect of sequence length on performance is weaker (Fig. A16). This is probably due to the fact that our environment has sparse obstacles and that larger frame-skip leads to effectively longer sequence and memory. Most results in our report use 4 policies for population based training and 8 environments per worker. These parameters also change the dynamics of learning. having more policies made agents with L=64 learn slower and L=8 learn slower. Further systematic investigation of these factors is required to establish broader applicability.
@@
Appendix B Implementation Details
B.1 Environment
@@
Navigation task does not require complex actions, thus we reduced the action space used in IMPALA (Espeholt et al., 2018), to facilitate training speed (Table A1).
Since we are mainly interested in modeling the way-finding aspect of navigation (Tolman, 1948; Kuipers, 2000), the MLP also receives average depth of the pixels, down-sampled to 10 horizontal pixels to aid motor control (going around walls and avoiding collisions). This procedure is supposed to mimic the sensory motor collision avoidance habits that do not depend on hippocampus. This depth information is directly routed to the decoder layer, so the representations in the recurrent core and its input are not directly influenced.
B.2 Multi-feature dynamics
As a more direct illustration for the recurrent dynamics of the sequence network $X_{t+1}=A\,X_{t}+B\,u_{t}$, we, here, provide the matrices $A$ and be for the multi-feature case:
$A=\begin{bmatrix}S&0&0&\dots&0\\ 0&S&0&\dots&0\\ \vdots&&\ddots&&\vdots\\ 0&\dots&&&S\end{bmatrix}_{F\ell\times F\ell},\qquad B=\begin{bmatrix}J&0&0&\dots&0\\ 0&J&0&\dots&0\\ \vdots&&\ddots&&\vdots\\ 0&\dots&&&J\end{bmatrix}_{F\ell\times F}\ .$ (5)
with
Note that the rows in $A$ corresponding to the first time step of each feature only contain 0s, i.e., sequences can only be started from DG input and not from the dynamics itself.
B.3 Architecture
Table A2: Network architecture of the CA3 agent.
Component Details
Encoder (ResNet) 3 conv blocks with residual layers, 16–32 channels, max pooling
MLP head Linear (3456 $\to$ 256), ReLU
DG projection Linear (256 $\to$ 16), BatchNorm (momentum 0.05, no affine), ReLU, intercept 2.43
[…]
Decoder MLP: Linear (1136 + 10 $\to$ 128), ReLU, Linear (128 $\to$ 128), ReLU
Critic Linear (128 $\to$ 1)
Actor Linear (128 $\to$ 5)
CA3 agent has learnable parameters in DG projection and from CA3 to Decoder, the CA3 output in the full version has 1136*128=149504 parameters. LSTM has learnable parameters $4*(m^{2}+mn)$ where m is hidden size and n is input and output size. Solving this gives a hidden size of roughly 137, which we used in the implementation. SSM agents have matching state size with CA3 agent. They also have block-diagonal recurrent weights, where the weights in each block are obtained through the original implementation by Gu et al. (2020) with zero-order-hold discretizations.
B.4 Actor Critic
We optimize an actor–critic objective with advantage estimates as in Espeholt et al. (2018), consisting of a clipped policy loss, a value regression loss, and an entropy bonus: